# Illuminating Blind Spots: Exploring LLMs as a Source of Targeted Synthetic Textual Data to Minimize High Confidence Misclassifications

### v2

@@
Abstract
Natural Language Processing (NLP) models optimized for predictive performance often make high confidence errors and suffer from vulnerability to adversarial and out-of-distribution data. Existing work has mainly focused on mitigation of such errors using either humans or an automated approach. In this study, we explore the usage of large language models (LLMs) for data augmentation as a potential solution to the issue of NLP models making wrong predictions with high confidence during classification tasks. We compare the effectiveness of synthetic data generated by LLMs with that of human data obtained via the same procedure. For mitigation, humans or LLMs provide natural language characterizations of high confidence misclassifications to generate synthetic data, which are then used to extend the training set. We conduct an extensive evaluation of our approach on three classification tasks and demonstrate its effectiveness in reducing the number of high confidence misclassifications present in the model, all while maintaining the same level of accuracy. Moreover, we find that the cost gap between humans and LLMs surpasses an order of magnitude, as LLMs attain human-like performance while being more scalable. [^1] Code is made available at: see supplementary materials.
Illuminating Blind Spots: Exploring LLMs as a Source of Targeted Synthetic Textual Data to Minimize High Confidence Misclassifications
Philip Lippmann, Matthijs T.J. Spaan, Jie Yang Delft University of Technology p.lippmann@tudelft.nl
1 Introduction
Natural Language Processing (NLP) models with the sole goal of optimizing predictive performance have achieved remarkable accuracy across a wide range of tasks but are often vulnerable to data that is adversarial in nature Papernot et al. (2016); Wang et al. (2019); Brown et al. (2020). Robustness – a model’s ability to deal with diverse and challenging data – plays a crucial role in addressing the vulnerability of NLP models to adversarial and out-of-distribution data Wang et al. (2022); Du et al. (2023), though it may be at odds with accuracy Tsipras et al. (2019).
[…]
Sentiment Prediction: Negative
Confidence: 0.98
[…]
Sentiment Prediction: Positive
Confidence: 0.94   Outcome: UU
[…]
Table 1: Rows 1 & 2 show an attack on a sentiment classifier using a perturbation method. Substitution of the perturbed text leads to change for a correct prediction to a wrong prediction at high confidence, leading to a UU. Row 3 is a generalized hypothesis that describes the cause of the UU. Row 4 is the exploration hypothesis based on the original hypothesis in Row 3.
@@ 1 Introduction
Figure 1: Procedure performed by human or LLM, beginning from a single UU present in a blind spot (denoted by a cross). This UU is then used to generate an initial hypothesis via abstraction. This abstraction hypothesis can then either by used to generate a synthetic sample that targets the existing blind spot or to generate a new hypothesis via exploration, which in turn is then used to generate a synthetic sample that targets an unknown blind spot.
Prediction errors can be categorized according to the model’s confidence in its prediction, with unknown unknowns (UUs) corresponding to misclassifications where the model is very confident Attenberg et al. (2015). UUs have been shown to cluster to form blind spots, which represent areas in the feature space that lead a model to produce high confidence misclassifications Lakkaraju et al. (2017); Liu et al. (2020). NLP models are prone to produce UUs as a consequence of text perturbations, such as the substitutions of relevant synonyms Ren et al. (2019). In the first two rows of table 1, we show an example of such a perturbation that successfully changes the predicted label, resulting in a UU. Discovering and mitigating blind spots is critical in high-stakes NLP tasks, where their presence can result in unwanted adverse outcomes, such as unreliable suicide prediction models Large et al. (2017) and biased sentencing decisions in criminal justice Crawford (2016).
The identification of UUs has been extensively explored in existing literature Attenberg et al. (2015); Bansal and Weld (2018); Vandenhof (2019); Liu et al. (2020); Han et al. (2021), with more recent researcher involving humans to aid blind spot identification efforts Cabrera et al. (2021); Han et al. (2021). However, current methods incorporate humans in a heuristic manner, not designing the task effectively to fully leverage their generalization abilities for blind spot mitigation. The non-triviality of mitigating blind spots in NLP models lies in the challenge of generalizing from previously identified individual UUs to the mitigation of unseen blind spots. It remains an open question how to best engage humans – or alternative agents, such as large language models (LLMs) – to effectively mitigate blind spots at scale.
To this end, we explore the characterization of existing blind spots by either humans or LLMs as a potential solution to the issue of NLP models making wrong predictions with high confidence, emphasizing the importance of generalization. Generalization refers to the idea that intelligent agents employ prior knowledge for new learning situations that share similarities with past experiences Gluck et al. (2011); Banich and Caccamise (2010); Allaway and McKeown (2020). We pose that by utilizing generalized patterns and similarities between seen and unseen high confidence misclassifications, blind spots can be effectively described by humans or LLMs. To accomplish this, we guide both humans and LLMs in formulating natural language hypotheses that describe the underlying principles contributing to the formation of blind spots. After generating hypotheses based on discovered UUs, we generate relevant additional synthetic samples which target UUs. In our study, we specifically compare the effectiveness of the synthetic data contributed by humans and LLMs, respectively.
We find that our approach is capable of greatly reducing the number of high confidence misclassifcations without decreasing accuracy. Further, we find LLMs exhibit a greater capacity for characterizing blind spots compared to humans, as evidenced by a larger average decrease in the number of UUs via our LLM-based approach (19.54%) compared to our human-based one (16.80%). Finally, the difference in cost per human-generated datum versus LLM-generated datum is substantial, with LLM-generated data being significantly more cost-effective.
2 Related Work
Before outlining our methodology, we introduce relevant prior research for approaches to high confidence misclassifications, as well as how others have tried to avoid such model behaviour.
2.1 Unknown Unknowns
Attenberg et al. (2015) introduce the concept of querying humans to find UUs in a game-like setting and show that there were patterns to the found UUs. Vandenhof (2019) proposes an approach to identify UUs where human-interpretable decision rules are learned to approximate how a model makes high-confidence predictions. Crowdworkers then contradict these rules by finding an instance that would classify as a UU. Han et al. (2021) propose an approach where crowdworkers continuously extend a dataset with additional UUs, that the chosen model is iteratively trained on. Cabrera et al. (2021) explore the use of crowdworkers to generate failure reports for NLP models to describe how or why the model failed. There are also algorithmic approaches to finding UUs, such as Lakkaraju et al. (2017), who propose utilizing an explore-exploit approach to find groups of UUs. Bansal and Weld (2018) extend this by proposing a utility model that rewards the degree to which the found UUs cover a sample distribution, thus encouraging the discovery of new blind spots. Instead, we do not find the UUs algorithmically, but instead use an LLM or crowdworkers to find existing UUs, extrapolate from these to unseen UUs, and generate synthetic data targeting these.
2.2 Model Calibration and Robust Training
The concept of UUs and blind spots is connected to model calibration (Guo et al., 2017; Minderer et al., 2021; Tian et al., 2023). A model that is well-calibrated will have its prediction confidence aligned with the likelihood of the correctness of the prediction. Investigating a model with blind spots is an examination of a model that is not well-calibrated. Therefore, our approach improves the calibration of the model.
In the case where the UUs are specifically generated through adversarial attacks, illumination of model blind spots is also related to robust training. UUs that populate these blind spots, when created by such attacks, may be identified as adversarial examples (Ribeiro et al., 2018; Wallace et al., 2019; Wang et al., 2020). This underscores the relationship between our proposed method and robust training practices, as well as the creation and identification of adversarial examples through adversarial attacks, with the aim of improving the robustness of the model (Madry et al., 2018; Pang et al., 2021). Unlike general robust training practices, our method focuses not on general robustness – but rather on high confidence misclassifications – and is not limited to just adversarial samples, as we also consider UUs that occur naturally, without any perturbation.
For the NLP research area, several approaches have been proposed to utilize synthetic data to expand training sets (Puri et al., 2020; Claveau et al., 2021). Synthetic textual data has previously been proposed specifically for sentiment classification tasks (Maqsud, 2015).  He et al. (2022) explore few-shot prompting LLMs to generate task specific synthetic training data. Unlike prior work, we propose a method to generate targeted synthetic data with the purpose of eliminating blind spots that lead to high confidence misclassifications.
3 Methodology
In our proposed approach, for the purpose of blind spot mitigation, a human or LLM is asked to perform one of three possible tasks: hypothesis generation via abstraction, hypothesis generation via exploration, or sample generation. These are described below and shown schematically in figure 1.
3.1 Problem Formulation
@@
For passive UU discovery, let the dataset be $\mathcal{D}=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$, where $x$ is the original text sample and $y$ the original label. Without having access to $y$, a predictive model $\theta$ is used to obtain $y=\theta(x)$ at a confidence $c\in[0,1]$. Formally, a UU occurs when (1) the wrong label is predicted by $\theta$ and (2) the prediction is made at $c\geq\tau$, indicating high confidence. In the case of proactive UU discovery, as is performed in this work, an additional technique is used to increase the misclassifications produced by $\theta$. For this, a black-box adversarial perturbation model $G$ generates perturbed samples $x^{\prime}=G(x)$, where $x^{\prime}\neq x$. Then $\theta$ is used to predict the new label for every sample $y^{\prime}=\theta(x^{\prime})$ at $c\in[0,1]$. The newly formed perturbed dataset $\mathcal{P}$ constitutes of these new labels and their corresponding $x^{\prime}$. A perturbation may be considered successful if $y^{\prime}\neq y$. If a perturbation occurs, there is an additional requirement for a UU: (3) $x^{\prime}$, regardless of its label indicated by the classifier, maintains the same underlying true label as $x$ post perturbation.
3.2 Generalization via Hypothesis Creation
For generalization, we present humans or LLMs with related pairs from $(x,y)\sim\mathcal{D}$ and $(x^{\prime},y^{\prime})\sim\mathcal{P}$ consisting of UUs and ask them to create hypotheses regarding the underlying causes of present blind spots. Due to the ability of humans to learn and generalize effectively using sparse amounts of data Lake et al. (2015), we can focus on subsets of UUs as relevant samples for the hypothesis generation task. In our study, hypotheses take the form of simple sentences that describe why a UU occurred and what information can be generalized from it to other potential UUs. As such, it does not act as a simple explanation of the failure case. Any hypothesis should cover multiple sentences indiscriminately, should the same characteristics persist across them. This allows a hypothesis to address multiple UUs that cluster together to form a blind spot and thus illuminate it. Examples of what these hypotheses might look like in practice are shown in the final two rows of table 1.
Abstraction Abstraction constitutes the usage of a single hypothesis to describe part of, or ideally, an entire blind spot, to expand the knowledge regarding that particular blind spot. During the abstraction step, we present UUs to both human participants and the LLM. They are shown a sample $(x_{i},y_{i})$ and its perturbed form $(x_{i}^{\prime},y_{i}^{\prime})$, and are tasked with generating a hypothesis that explains any factors that could have led to it being a UU. We frame the task in a way that this could be related to a pattern, semantics, syntax, specific words, or something else between the original and perturbed samples. They are tasked with employing abstract reasoning to come up with a natural language hypothesis, while being reminded that this should generalize to other, similar UUs. An example of abstraction being performed is shown in table 3.
Exploration Exploration as a strategy goes beyond attempting to reason about blind spots that result from discovered UUs. Instead, exploration focuses on using existing hypotheses to generalize to a new hypothesis and thus find additional, previously undiscovered blind spots to gain new knowledge regarding the model’s failure modes. Humans and LLMs are encouraged to use extrapolative thinking to form a hypothesis that they believe covers new potential failure modes. To this end, we request they ensure that the generated hypothesis is dissimilar compared to the already established hypothesis while targeting another potential weakness. For the exploration step, we present human-generated hypotheses to the human participants and vice versa. An example of exploration being performed is shown in table 4.
3.3 Synthetic Sample Generation
After generating hypotheses via abstraction or exploration, participants, whether human or LLM, are asked to generate synthetic samples, which we denote as $\bar{x}^{\prime}$, that follow these new hypotheses. These new samples must fit the overall structure of the used dataset, as well as be applicable to the overall context of the task (e.g., a movie-related sample for a movie review dataset). We thus obtain additional samples that should, according to our intuition, correspond to blind spots that the predictive model $\theta$ might have. Using these samples the training dataset is extended, resulting in $\mathcal{E}$, which consists of $\mathcal{D}$ and the additional generated samples $\bar{x}^{\prime}$ and their corresponding labels $\bar{y}^{\prime}$, which we also ask participants to provide. The generation procedure is the same, regardless of whether the used hypothesis was obtained via abstraction or exploration, and humans perform generation exclusively on human-generated hypotheses and vice versa.

### v3

# Illuminating Blind Spots of Language Models with Targeted Agent-in-the-Loop Synthetic Data

@@
Abstract.
Language models (LMs) have achieved impressive accuracy across a variety of tasks but remain vulnerable to high-confidence misclassifications, also referred to as unknown unknowns (UUs). These UUs cluster into blind spots in the feature space, leading to significant risks in high-stakes applications. This is particularly relevant for smaller, lightweight LMs that are more susceptible to such errors. While the identification of UUs has been extensively studied, their mitigation remains an open challenge, including how to use identified UUs to eliminate unseen blind spots. In this work, we propose a novel approach to address blind spot mitigation through the use of intelligent agents – either humans or large LMs – as teachers to characterize UU-type errors. By leveraging the generalization capabilities of intelligent agents, we identify patterns in high-confidence misclassifications and use them to generate targeted synthetic samples to improve model robustness and reduce blind spots. We conduct an extensive evaluation of our method on three classification tasks and demonstrate its effectiveness in reducing the number of UUs, all while maintaining a similar level of accuracy. We find that the effectiveness of human computation has a high ceiling but is highly dependent on familiarity with the underlying task. Moreover, the cost gap between humans and LMs surpasses an order of magnitude, as LMs attain human-like generalization and generation performance while being more scalable.
Unknown Unknowns, Blind Spots, Language Models, Model Robustness, Human-in-the-loop
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing User studiesccs: Computing methodologies Natural language generationccs: Computing methodologies Learning settings
1. Introduction
@@
Figure 1. In a sentiment classification task, we begin with a UU resulting from a perturbation – denoted by a cross in the feature space. This UU is then used to generate an initial hypothesis via abstraction through human computation or an LM. This abstraction hypothesis can then either by used to generate a synthetic samples that target the existing blind spot or to generate a new hypothesis via extrapolation, which in turn is then used to generate synthetic samples targeting an unseen blind spot.
Language models (LMs) have achieved remarkable accuracy across a wide range of predictive tasks, but remain vulnerable to out-of-distribution data (Papernot et al., 2016; Wang et al., 2019; Brown et al., 2020a). Small, lightweight LMs – while easier to train and run on limited hardware, and therefore favored in domain-specific applications – are especially prone to UUs due to their reduced robustness (Wang et al., 2022; Du et al., 2023). Larger LMs, although generally more robust, require significant computational resources for both training and inference, limiting their usability (Touvron et al., 2023). This vulnerability often leads to prediction errors, including in high-stakes applications such as suicide prevention (Large et al., 2017) and criminal justice sentencing (Crawford, 2016), where reliable and unbiased predictions are critical. A particularly challenging class of errors, referred to as unknown unknowns (UUs), occurs when the model confidently misclassifies an input as the incorrect label (Attenberg et al., 2015). These UUs tend to cluster into blind spots in the feature space, areas where the model consistently produces high-confidence misclassifications due to biases in the training data (Lakkaraju et al., 2017; Liu et al., 2020). On the left side of figure 1 we show an example of a mispredicted label at a high confidence, resulting in a UU, that forms part of a blind spot.
The identification of UUs and blind spots has been extensively studied (Attenberg et al., 2015; Bansal and Weld, 2018; Vandenhof, 2019; Liu et al., 2020), including approaches involving human oversight to aid in detection (Cabrera et al., 2021; Han et al., 2021). Mitigating blind spots – especially how to move from identified blind spots to unseen ones – remains an unresolved challenge. Simple approaches to tackling only already discovered blind spots, such as relabeling previously identified UUs and using them for additional training (Han et al., 2021), do not scale and fall short of ensuring a holistic reduction in blind spots. Thus the only blind spots of the model that can be illuminated using such reactive approaches are those that correspond to seen data, with those that correspond to unseen data remaining out of reach.
In this paper, we introduce an agent-in-the-loop workflow that proactively mitigates blind spots of LMs by employing intelligent agents – either humans or large LMs – to characterize blind spots and subsequently generate targeted synthetic data. We pose that the key to mitigating these blind spots lies in the generalization abilities of the agent, allowing them to hypothesize patterns of discovered UUs and similarities between seen and unseen UUs using prior knowledge (Gluck et al., 2011; Banich and Caccamise, 2010; Allaway and McKeown, 2020). To this end, we guide agents to formulate these hypotheses in natural language, either describing the found blind spot consisting of discovered UUs (abstraction) or reasoning about undiscovered blind spots (extrapolation), as is shown in figure 1. Using these hypotheses, we guide agents toward the generation of synthetic samples targeted at blind spots, improving the robustness of LMs through subsequent retraining by reducing the number of high-confidence misclassifications without sacrificing overall predictive accuracy. Our workflow is designed to flexibly integrate intelligence from both humans and LMs, with specific mechanisms to incorporate human computation or LMs. Additionally, the workflow can incorporate existing adversarial attack methods to proactively illuminate blind spots, further enhancing its adaptability and effectiveness.
Our workflow proves to be a viable means of distilling knowledge from intelligent agents to small LMs, making them more robust while maintaining their lightweight advantages. Through our comprehensive experiments, we find that our method is capable of substantially reducing the number of high-confidence misclassifcations without decreasing accuracy. On average, we are able to reduce the number of UUs by 19.08%. Further, we show that for our method LMs are more effective overall than human agents, achieving a 22.37% reduction in UUs compared to a 15.78% reduction when using human-generated data. Additionally, LM-generated data are far more economical, making them a more scalable solution for improving the robustness of small models. Finally, we observe that humans surpass LMs in certain tasks, particularly those that align more closely with human intuition due to their greater familiarity to participants.
In summary, the contributions of this paper are as follows:
  • A new workflow that utilizes the generalization capabilities of intelligent agents to mitigate blind spots, employing targeted synthetic data generation through an identify-characterize-generate approach.
  • A comparative study on the efficacy of humans and LMs in applying our workflow to a variety of classification use cases and classification models, demonstrating the task-dependency of human contributions and scalability of LM-derived data.
2. Agent-in-the-Loop Targeted Data Generation
Our proposed approach to blind spot mitigation involves engaging a human or LM in three tasks: hypothesis generation via abstraction, hypothesis generation via extrapolation, and synthetic sample generation. These tasks are designed to characterize and mitigate blind spots, ultimately reducing high-confidence misclassification. The workflow is schematically illustrated in figure 1. The human computation component of our study is implemented through a survey study, the details of which are provided in appendix B, while the equivalent LM prompts are given in appendix C.
2.1. Problem Formulation
For UU discovery, let the dataset be $\mathcal{D}=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$, where $x$ is the original text sample and $y$ the original ground truth label. Without having access to $y$, a predictive model $\theta$ is tasked with generating a label prediction $y_{p}=\theta(x)$ at a confidence $c\in[0,1]$. Formally, a UU occurs when (1) $\theta$ predicts the wrong label $y_{p}\neq y$ and (2) the prediction is made with high confidence $c\geq\tau$.
@@ 2.1. Problem Formulation
In this work, in addition to dealing with the blind spots that naturally occur in models as a result of training, we make use of adversarial UU discovery, where we increase the number of misclassifications by introducing perturbations. For this, a black-box adversarial perturbation model $G$ generates perturbed samples $\bar{x}=G(x)$, where $\bar{x}\neq x$. The model $\theta$ is then used to predict new labels $y^{\prime}_{i}=\theta(\bar{x}_{i})$ at a confidence $c$. The resulting perturbed dataset, denoted $\mathcal{P}$, consists of the new samples and predicted labels $(\bar{x},y^{\prime})$. If a perturbation occurs, there is an additional requirement for a misclassification to be considered a UU: (3) $\bar{x}$, regardless of its label indicated by $\theta$, maintains the same underlying true label $y$ as $x$ post perturbation.
Given a predictive model $\theta$ trained on a dataset $\mathcal{D}$, our objective is to mitigate UUs produced by $\theta$. To systematically reduce high-confidence misclassification, we seek to identify patterns in discovered UUs and generate targeted synthetic data $\{x^{s},y^{s}\}$ for a set of UUs, where $x^{s}$ is the synthetic label and $y^{s}$ represents the corresponding ground truth label for the synthetic sample. This data is then used to further train $\theta$ and thus reduce the blind spots present.
2.2. Generalization via Hypothesis Creation
For UU mitigation, we employ intelligent agents (humans or large LMs) to generalize from identified UUs to create hypotheses in natural language regarding the underlying causes of these UUs. As we use perturbations, such hypotheses are based on pairs of original and perturbed samples, $(x_{i},y_{i})\sim\mathcal{D}$ and $(\bar{x}_{i},y_{i}^{\prime})\sim\mathcal{P}$. Humans are adept at using sparse data to generalize (Lake et al., 2015), and this task exploits that capability by focusing on subsets of UUs. Each hypothesis describes the shared characteristics that explain why certain UUs occur and how these characteristics might generalize to other, unseen UUs. The goal is not merely to explain individual failure cases but to construct hypotheses that address multiple UUs clustering together into a blind spot. In doing so, we can illuminate patterns within the feature space that the model is consistently misclassifying. To this end, we pursue two distinct but complementary strategies: abstraction and extrapolation.
2.2.1. Abstraction
Abstraction involves generating a hypothesis on why a specific UU occurred that generalizes across a set of closely related UUs, revealing underlying patterns within a blind spot. In this step, the intelligent agent is provided with an original sample $(x_{i},y_{i})$ and, if adversarial perturbations are used, its perturbed counterpart $(\bar{x}_{i},y_{i}^{\prime})$. Then the agent is tasked with reasoning abstractly about the factors leading to this UU. Specifically, we instruct them to consider whether these factors involve semantics, syntax, specific words, or something else in the samples that could be the cause of the high-confidence misclassification. This is to guide the agent to identify what most likely contributes to the UU without prescribing rigid criteria, leaving room for creative thinking and allowing the agent to explore unforeseen or nuanced factors. The hypothesis is in natural language and should generalize across other UUs that share these characteristics, expanding our understanding of the particular blind spot the UU corresponds to. Compared to a mitigation approach that only makes use of a simple reactive relabeling of found UUs, our method comes with the additional advantage that it builds up a corpus of human-interpretable error reports on seen errors of the classification model. An example of hypothesis generation via abstraction is shown in figure 2.
2.2.2. Extrapolation
Extrapolation extends the process of hypothesis creation beyond trying to describe discovered blind spots, encouraging the agent to use existing hypotheses and sample pairs (used during abstraction) to uncover new blind spots. This task emphasizes extrapolation, asking the agent to hypothesize new failure modes – also in natural language – that differ from those previously identified. Extrapolative thinking has previously been shown to be a human strong suit (Bartlett, 1958). By ensuring that the new hypotheses are dissimilar from those used for abstraction, we aim to discover new regions in the feature space where the model may be prone to high-confidence misclassification. To avoid the agent overextrapolating, we specifically instruct them to focus on the same topic but reason if a different possible factor from semantics, syntax, specific words could be at fault that was not mentioned in the abstraction hypothesis. In this step, we present only human-generated hypotheses to human participants and vice versa. An example of extrapolation is shown in figure 2.
2.3. Synthetic Sample Generation
Once hypotheses have been generated via abstraction or extrapolation, the agent is tasked with generating synthetic samples. These synthetic samples must align with the structure and context of the original dataset while reflecting the characteristics of the generated hypotheses. For instance, if the dataset consists of movie reviews, the synthetic samples should maintain the form and tone of movie review-related text. The goal of this step is to create new data points that correspond to the blind spots identified during hypothesis generation. These synthetic samples are added to the training dataset, resulting in a dataset that is extended for each synthetic sample and its corresponding label $\mathcal{E}=\mathcal{D}\cup\{x^{s}_{i},y^{s}_{i}\}$, where the label is provided by the agent. By incorporating these new samples into training, we aim to enhance the robustness of the predictive model $\theta$ by reducing its susceptibility to high-confidence misclassifications. The sample generation process is uniform, regardless of whether the hypothesis was obtained through abstraction or extrapolation. Humans generate samples based on human-created hypotheses, and LMs do the same for LM-generated hypotheses. An example of this type of sample generation from human and LM agents for abstraction and extrapolation is shown in figure 2.
Example of full workflow for a single UU sample using human computation and LMs on the IMDB Dataset Original Sample: Footage of busy highways, planes flying overhead, poor lightning effects, soldiers doing their thing, they’re all used in a [[haphazard]] manner, sometimes repeatedly, and they often only have a loose connection to the story Wood is trying to tell. […] Perturbed Sample: Footage of busy highways, planes flying overhead, poor lightning effects, soldiers doing their thing, they’re all used in a [[thoughtless]] manner, sometimes repeatedly, and they often only have a loose connection to the story Wood is trying to tell. […]   Human Abstraction Hypothesis: “Haphazard can be intentional, vs thoughtless is unplanned. It feels more of an unwillingness rather than a lack of ability.” LM Abstraction Hypothesis: “The replacement of specific adjectives in the perturbed sample contributes to misclassification. The adjective ”thoughtless” in the perturbed sample, may change the perception of the director’s skills and use of stock footage. This alteration in sentiment could lead to a different classification outcome.”   Human Extrapolation Hypothesis: “The hypothesis could be that the words were potentially swapped or replaced, such as ”bad” being replaced by ”evil” which could lead to an incorrect prediction as it is similar but could have a slightly different meaning in some contexts.” LM Extrapolation Hypothesis: “The use of the word ”hypocrite” may have caused the misclassification. The word ”hypocrite” implies that the person is saying one thing but doing another, which may be considered a negative trait. However, some people may not interpret the person in the sample as a hypocrite, leading to a difference in sentiment analysis. This difference in interpretation may have caused the misclassification by the algorithm.”   Human Abstraction Sample: “The scenes felt like they were slapped together with no care at all. It’s not that the director didn’t have the skill they just didn’t bother to try harder. Haphazard is at least creative but thoughtless just shows they didn’t wanna put the effort in.” LM Abstraction Sample: “Skyward Dreams had potential, but the director’s use of stock footage feels thoughtless. Scenes of bustling cities and planes flying overhead are inserted without any real purpose, disrupting the flow of the narrative. […]” Human Extrapolation Sample: “The CGI in was straight up evil. The way the effects looked completely ruined the immersion for me, and it felt like the creators didn’t even care about quality. I get that sometimes budget is an issue, but this was just on another level. […]” LM Extrapolation Sample: “The protagonist of The Final Betrayal is a true hypocrite. Throughout the film, he preaches loyalty and honesty to his friends, yet secretly manipulates and betrays them behind their backs. This hypocrisy is central to the film’s conflict, as the character’s outward morality sharply contrasts with his deceitful actions. Despite this glaring flaw, some viewers may interpret his behavior as a survival tactic in a harsh world, rather than outright hypocrisy. […]”
Figure 2. Example of hypothesis generalization using abstraction for the IMDB dataset. The abstraction is performed by a human or LLM based on original and perturbed samples.